Awolflover1 (awolflover1):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 5) and (5, 1) C. (10, 6) and (5, 1) D. (10, 6) and (1, 5)

OpenStudy (anonymous):

The answer is C because y = x - 4 6 = 10 - 4 6 = 6 y = x - 4 1 = 5 - 4 1 = 1